OpenStudy (lexilove262):

Inverse of functions help F (×)1-4 x I desperate please help. I thought I understood in class but now I don't

OpenStudy (lexilove262):

It supposed to be f (x)10-4x

satellite73 (satellite73):

\[f(x)=10- x\]??

satellite73 (satellite73):

or maybe \[f(x)=10-4x\]

OpenStudy (lexilove262):

I got f (x) =10-×/4

satellite73 (satellite73):

what is the original function ?

OpenStudy (lexilove262):

But I wasn't sure if I was right or wrong

OpenStudy (lexilove262):

We didn't have original functions just the function I first posted f (x)=10-4x

satellite73 (satellite73):

ok then if you got \[f^{-1}(x)=\frac{10-x}{4}\] you are right

OpenStudy (lexilove262):

Okay that's what I got I just needed someone to make sure because I did the whole paper like that thanks☺

satellite73 (satellite73):

yw